OpenStudy (anonymous):

how do i determine if a function is exponential or logarithmic if im only given the domain and range

OpenStudy (anonymous):

exponents have an infinite domain and finite range logarithms have a finite domain and infinite range

OpenStudy (anonymous):

it gave me D={-2, -2, 0, 1, 2} R={0.25, 1, 4, 16, 64}

OpenStudy (anonymous):

This is not a function as -2 must map to two different values in the range and functions only have one output for any given input

OpenStudy (anonymous):

the problem just shows that and asks whether it logarithmic or exponential

OpenStudy (anonymous):

If you meant -1 then this is exponential, if you plot the data it starts to increase slowly and then speeds up

OpenStudy (anonymous):

what about D={0.25, 1, 4, 16, 64} R={-2, -1, 0, 1, 2}

OpenStudy (anonymous):

(although all logarithmic functions are, are simply the inverse of an exponential function, so technically all logarithmic functions are exponential as well.

OpenStudy (anonymous):

in that one they changed the domain and range so the first was exponential, this one is the inverse so it must be logarithmic.

OpenStudy (anonymous):

and how can i tell if a graph is logarithmic or exponential if im only given a graph, nothing else

OpenStudy (anonymous):

If it starts off flat and then curves up (or down) it's exponential. If it starts off going up (or down) and flattens out it's logarithmic.

OpenStudy (anonymous):

oh! ok, thank you :) do you know anything about equations containing radicals and exponents by any chance?

OpenStudy (anonymous):

Ohhh, I know a thing or two about those guys. :)

OpenStudy (anonymous):

how would i solve 1/4^3x=2

OpenStudy (anonymous):

Is it:\[(1/4)^{3x}=2\]

OpenStudy (anonymous):

the x isnt attatched to the exponent,

OpenStudy (anonymous):

ok so: \[(1/4)^{3}x=2\]

OpenStudy (anonymous):

yes, and the fraction isnt in parenthesis

OpenStudy (anonymous):

First raise 1/4 to the third power: 1/4^3 = 1/64 (4^3 = 4*4*4) so 1/64x = 2 multiply both sides by 64 x = 128

OpenStudy (anonymous):

wait i made a mistake, its 1/4x^3=2

OpenStudy (anonymous):

oh, that makes a little bit of a difference... First multiply both sides by 4 x^3 = 8 next raise both sides to the (1/3) power x = (8)^(1/3) x = 2

OpenStudy (anonymous):

wht if i had 1/8x^6-3=5

OpenStudy (anonymous):

Working backwards get rid of the -3 first 1/8x^6 = 8 next multiply by 8 x^6 = 64 now raise both sides to the 1/6 power x = (64)^(1/6) x = 2

OpenStudy (anonymous):

One way to get rid of exponents is to raise both sides to 1/n where n is the exponent.

OpenStudy (anonymous):

how come i have to make the exponent 1/6?

OpenStudy (anonymous):

oh! ok, but how does it equal to 2

OpenStudy (anonymous):

if you have a calculator you can type in 64^(1/6) and it should give you the answer of 2. Or you can ask yourself what number times itself 6 times is 64? And then use guess and check to find 2. (I like the calculator way myself)

OpenStudy (anonymous):

oh i see, what do i do if i have (x-1)^2/3-13=3 i know i have to add 13 to both sides but what do i do with the exponent